# LLM4Cell: A Survey of Large Language and Agentic Models for Single-Cell Biology

### v2

@@ 1 Introduction
Large language models (LLMs) are transforming biomedical discovery by linking molecular patterns with knowledge encoded in text.  (Zhang et al., 2025a; Aljohani et al., 2025; Bi et al., 2024). Prior work on computational workflows and biological prediction tasks demonstrates how learned representations capture complex molecular relationships and enhance practical performance.  (Lam et al., 2024; Bhattacharya et al., 2024; Dip et al., 2024; Nam et al., 2024; Shuvo et al., 2024) In single-cell biology, where each experiment measures thousands of genes across millions of cells, LLMs promise to unify gene expression, chromatin accessibility, and spatial organization under a shared, interpretable framework. Early foundation models such as scGPT  (Cui et al., 2024), Geneformer  (Theodoris et al., 2023), and scFoundation  (Hao et al., 2024) learn transferable representations directly from single-cell data, while emerging systems like CellLM  (Zhao et al., 2023), scAgent  (Mao et al., 2025), and CellVerse  (Zhang et al., 2025b) extend these capabilities toward reasoning, dialogue, and autonomous analysis. Together, they signal a shift from statistical inference to language-driven, interpretable, and generalizable cell intelligence.
@@ 2 Related Work
The intersection of large language models (LLMs) and single-cell biology has gained rapid momentum, motivating several early surveys and benchmarks. The ACL 2025 survey on foundation models for single-cell biology (Zhang et al., 2025a) outlines pretrained and fine-tuned architectures across protein, gene, and cell representations. Broader reviews such as Lan et al. (2025) and Bian et al. (2024b) discuss “large cellular models” (LCMs) and trends in scaling, tokenization, and transfer learning across biological data. Benchmark studies including Single-Cell Omics Arena (Liu et al., 2024) and CellVerse (Zhang et al., 2025b) evaluate LLMs on annotation and question answering, while works like scInterpreter (Guo et al., 2023) and scExtract (Wu and Tang, 2025) demonstrate practical integration into annotation and curation pipelines.
@@ 3 Datasets
Transcriptomic atlases such as Tabula Sapiens, Tabula Muris  (Consortium* et al., 2022), and the Human Cell Atlas  (Travaglini et al., 2020) remain dominant for foundation-model training and ontology-aware annotation. Chromatin-accessibility data (e.g. Cusanovich mouse atlas (Cusanovich et al., 2018), human adult/fetal scATAC (Domcke et al., 2020)) map regulatory states but remain sparse and heterogeneous. Multi-omic resources (e.g. TEA-seq  (Swanson et al., 2021), DOGMA-seq  (Mimitou et al., 2021), CITE-seq (Stoeckius et al., 2017)) link RNA, ATAC, and protein modalities, providing supervision for cross-view alignment. Spatial technologies (e.g. Visium  (Oliveira et al., 2025), Slide-seqV2  (Stickels et al., 2021), MERFISH  (Chen et al., 2015), Stereo-seq  (Chen et al., 2022)) connect molecular profiles to tissue architecture, while functional datasets (e.g. Perturb-seq (Replogle et al., 2022)) benchmark causal reasoning. Plant resources (e.g. scPlantDB  He et al. (2024), Arabidopsis E-CURD-4  (Shulse et al., 2019)) extend modeling beyond animal systems, opening cross-kingdom evaluation.
@@ 4.1 Foundation Models
Foundation models learn transferable cell and gene embeddings directly from large-scale scRNA-seq without explicit labels. Representative systems such as scGPT  (Cui et al., 2024), Geneformer  (Theodoris et al., 2023), and scFoundation  (Hao et al., 2024) pretrain on multi-tissue atlases exceeding one million cells, using masked-gene or rank-based reconstruction to capture expression context. Variants like tGPT  (Shen et al., 2023), scBERT  (Yang et al., 2022), and scGraphformer  (Fan et al., 2024) treat genes as tokens or graph nodes, while cross-species models (UCE  (Rosen et al., 2023), GeneCompass (Yang et al., 2024b)) apply contrastive alignment across homologous genes. These models underpin most single-cell LLM pipelines, offering strong transfer for annotation and integration but limited ontology grounding and explainability, motivating later text-bridged and reasoning frameworks.
@@ 4.2 Text-Bridge LLMs
Text-Bridge models couple molecular embeddings with biomedical language to ground single-cell representations in semantics and ontology. Systems such as scELMo  (Liu et al., 2023), CellLM  (Zhao et al., 2023), and GenePT  (Chen and Zou, 2024) align gene or cell embeddings with textual descriptors, while Cell2Sentence  (Levine et al., 2024) and Cell2Text  (Kharouiche et al., 2025) translate expression profiles into natural-language summaries. Using dual-encoder or encoder–decoder architectures, these frameworks employ contrastive or prompt-based alignment between molecular and text encoders (e.g., BioBERT). They enhance interpretability and enable zero-shot annotation but depend heavily on curated corpora and remain non-agentic, serving as a bridge between foundation and reasoning frameworks.
@@ 4.3 Spatial and Multimodal Models
Spatial and multimodal frameworks integrate gene expression with spatial coordinates, histology, or additional omics to capture tissue architecture. Representative models include TransformerST  (Lu et al., 2024), spaLLM (Ji et al., 2024), and OmiCLIP (Cui et al., 2025) for spatial mapping, and scMMGPT (Shi et al., 2025) and FmH2ST for RNA–ATAC–protein integration. They use multi-branch Transformers with modality-specific encoders and cross-attention fusion, trained on datasets such as Visium DLPFC and MERFISH. These models achieve strong spatial alignment and biological realism but face heterogeneous resolutions, limited open benchmarks, and high computational cost.
@@ 4.4 Epigenomic Models
Epigenomic foundation models extend LLM concepts to chromatin-accessibility and regulatory data such as scATAC-seq. EpiFoundation  (Wu et al., 2025), EpiBERT (Javed et al., 2025), and EpiAttend (Li et al., 2022) learn cis-regulatory patterns from ENCODE and other compendia, while GeneMamba and scMamba use efficient state-space layers for long-range dependency modeling. Cross-omic variants (ChromFound (Jiao et al., 2025), GET (Fu et al., 2025)) jointly embed RNA and ATAC to infer gene-regulatory networks. These models improve biological grounding but remain constrained by sparse data and lack unified benchmarks across regulatory modalities.
@@ 4.5 Agentic Frameworks
Agentic systems integrate pretrained models with reasoning modules for autonomous single-cell analysis. Frameworks such as scAgent (Mao et al., 2025), CellVerse (Zhang et al., 2025b), and Teddy (Chevalier et al., 2025) combine domain-specific encoders with LLM controllers that plan tasks, query ontologies, and interface with tools or APIs. EpiAgent (Chen et al., 2025b) extends this paradigm to regulatory genomics. These systems enable dialogue-based annotation and multi-step reasoning but lack standardized benchmarks for reasoning fidelity and rely on underlying LLM reliability.
@@ 5.1 Annotation and Ontology Mapping
Annotation is the most common single-cell task, spanning automated cell-type labeling, ontology alignment, and cross-dataset harmonization. Foundation models such as scGPT, scFoundation, and scBERT achieve high accuracy on atlases like Tabula Sapiens and the Human Cell Atlas. Text-bridge models (CellLM  (Zhao et al., 2023), scELMo  (Liu et al., 2023), GenePT  (Chen and Zou, 2024)) add ontology-based interpretability, and agentic systems (scAgent) perform multi-step annotation reasoning via LLM controllers. Models are fine-tuned on curated references (e.g., Azimuth PBMC, HCA Lung) using masked-gene or contrastive objectives. Evaluation uses label accuracy or ARI; key limitations include rare-cell detection, cross-species consistency, and lack of reasoning benchmarks.
@@ 5.2 Trajectory and Perturbation Modeling
Trajectory modeling captures dynamic state transitions and causal responses to interventions. Geneformer  (Theodoris et al., 2023), scGPT (Cui et al., 2024), and scRDiT  (Dong et al., 2025) model temporal or perturbation trajectories using denoising and contrastive objectives, trained on datasets such as Replogle 2022 Perturb-seq and sci-Plex. Epigenomic models (EpiFoundation (Wu et al., 2025) ,EpiAgent (Chen et al., 2025b)) extend to regulatory responses through chromatin context and ontology reasoning. Metrics include expression-profile correlation and target-gene recovery. While these models predict single perturbations effectively, performance degrades for combinatorial or long-range effects, underscoring the need for unified causal benchmarks.
@@ 5.3 Multi-omic Integration
Multi-omic integration seeks unified representations across paired RNA, ATAC, and protein modalities to capture gene-regulatory and signaling relationships. Representative models include scMMGPT (Shi et al., 2025), GET (General Expression Transformer) (Fu et al., 2025), ChromFound (Jiao et al., 2025), and epigenomic-aware variants such as EpiFoundation (Wu et al., 2025) and EpiBERT (Javed et al., 2025). These systems train on joint datasets like 10x Multiome PBMC 10k, TEA-seq, DOGMA-seq, and ASAP-seq, using cross-modal transformers or contrastive alignment between expression and accessibility features.
@@ 5.4 Spatial Mapping and Deconvolution
Spatial mapping links molecular profiles to their tissue locations, enabling inference of cell organization and microenvironmental context. Representative models include TransformerST  (Lu et al., 2024), spaLLM , OmiCLIP, FmH2ST, HEIST  (Madhu et al., 2025), and Spatial2Sentence (Chen et al., 2025a), trained on datasets such as Visium DLPFC, Slide-seqV2, MERFISH, and Xenium. They combine spatial coordinates, histology features, and gene-expression embeddings through cross-attention or contrastive alignment to reconstruct cell or spot-level maps.
@@ 5.5 Regulatory Network and Pathway Inference
Regulatory-network and pathway inference focuses on identifying gene–gene or enhancer–promoter dependencies that drive transcriptional programs. Representative models include GeneMamba (Qi et al., 2025a), scMamba (Yuan et al., 2025), EpiFoundation (Wu et al., 2025), EpiBERT (Javed et al., 2025), ChromFound (Jiao et al., 2025), and GET (Fu et al., 2025), trained on large-scale ATAC and multiome atlases such as ENCODE, human cCRE datasets, and 10x Multiome PBMC. These models treat chromatin regions or genes as tokens and learn context-dependent regulatory relationships through self- or cross-attention.
GeneMamba (Qi et al., 2025a) and scMamba employ state-space layers for efficient modeling of long-range chromatin dependencies, while EpiBERT and EpiFoundation use masked-region reconstruction to capture enhancer–promoter coupling. ChromFound and GET link RNA and accessibility signals to infer transcription-factor activity and causal directionality. Evaluation typically uses motif enrichment, AUROC for known interactions, or pathway-level overlap with KEGG or Reactome references. These models improve biological grounding and causal interpretability but remain constrained by sparse training data and lack of unified ground-truth regulatory benchmarks.
@@ 5.6 Cross-Species Translation
Cross-species translation aims to transfer cell and gene representations between organisms, enabling comparative and evolutionary analysis of single-cell atlases. Representative models include iSEEEK  (Shen et al., 2022), UCE (Universal Cell Embeddings)  (Rosen et al., 2023), GeneCompass  (Yang et al., 2024b), and scPlantLLM  (Cao et al., 2025), trained on paired or homolog-mapped datasets such as tthe Mouse Cell Atlas, Tabula Muris, and scPlantDB. These systems learn alignment between orthologous genes or conserved cell states to support zero-shot annotation across species.
iSEEEK  (Shen et al., 2022) integrates over 11 million human and mouse cells using shared gene-token vocabularies, while UCE  (Shen et al., 2023) applies masked autoencoding and contrastive alignment across ortholog groups. GeneCompass  (Yang et al., 2024b) extends this with knowledge-informed pretraining that links species via ontology-derived embeddings, and scPlantLLM adapts the paradigm to plant single-cell data across 17 species. Evaluation uses transfer accuracy and embedding consistency across species pairs. Despite improved cross-domain generalization, these models remain limited by incomplete ortholog mappings and domain shifts in sequencing depth and tissue composition.
@@ 5.7 Generation and Simulation
Generative modeling aims to synthesize realistic single-cell profiles, simulate perturbations, or reconstruct missing modalities. Representative approaches include scGPT, scFoundation, CellFM  (Zeng et al., 2025), and Geneformer, which employ decoder or diffusion-style architectures to model the probability distribution of gene-expression states. These models are pretrained on large atlases such as Tabula Sapiens and HCA, then fine-tuned to generate new cell embeddings or perturbation outcomes.
@@ 5.8 Drug-Response Prediction
Drug-response prediction models aim to infer cellular transcriptional outcomes following chemical or genetic perturbations, supporting virtual screening and mechanism discovery. Representative frameworks include Geneformer (Theodoris et al., 2023), scGPT  (Cui et al., 2024), EpiFoundation (Wu et al., 2025), and EpiAgent (Chen et al., 2025b), trained or evaluated on large-scale pharmacotranscriptomic resources such as sci-Plex, Replogle 2022 Perturb-seq, and the ARC Virtual Cell Challenge. These systems learn conditional embeddings that map control to perturbed states, enabling zero-shot prediction of unseen compounds or targets.

### v3

@@ 1 Introduction
Large language models (LLMs) are transforming biomedical discovery by linking molecular patterns with knowledge encoded in text.  (zhang2025survey; aljohani2025comprehensive; bi2024ai). Prior work on computational workflows and biological prediction tasks demonstrates how learned representations capture complex molecular relationships and enhance practical performance.  (lam2024large; bhattacharya2024large; dip2024deepage; nam2024using; shuvo2024assessing) In single-cell biology, where each experiment measures thousands of genes across millions of cells, LLMs promise to unify gene expression, chromatin accessibility, and spatial organization under a shared, interpretable framework. Early foundation models such as scGPT  (cui2024scgpt), Geneformer  (theodoris2023transfer), and scFoundation  (hao2024large) learn transferable representations directly from single-cell data, while emerging systems like CellLM  (zhao2023large), scAgent  (mao2025scagent), and CellVerse  (zhang2025cellverse) extend these capabilities toward reasoning, dialogue, and autonomous analysis. Together, they signal a shift from statistical inference to language-driven, interpretable, and generalizable cell intelligence.
@@ 2 Related Work
The intersection of large language models (LLMs) and single-cell biology has gained rapid momentum, motivating several early surveys and benchmarks. The ACL 2025 survey on foundation models for single-cell biology (zhang2025survey) outlines pretrained and fine-tuned architectures across protein, gene, and cell representations. Broader reviews such as lan2025large and bian2024general discuss “large cellular models” (LCMs) and trends in scaling, tokenization, and transfer learning across biological data. Benchmark studies including Single-Cell Omics Arena (liu2024single) and CellVerse (zhang2025cellverse) evaluate LLMs on annotation and question answering, while works like scInterpreter (guo2023scinterpreter) and scExtract (wu2025scextract) demonstrate practical integration into annotation and curation pipelines.
@@ 3 Datasets
Transcriptomic atlases such as Tabula Sapiens, Tabula Muris  (tabula_sapiens2022), and the Human Cell Atlas  (travaglini2020_hlca) remain dominant for foundation-model training and ontology-aware annotation. Chromatin-accessibility data (e.g. Cusanovich mouse atlas (cusanovich2018_science), human adult/fetal scATAC (zhang2021_encode)) map regulatory states but remain sparse and heterogeneous. Multi-omic resources (e.g. TEA-seq  (swanson2021_natbiotech), DOGMA-seq  (mimitou2021scalable), CITE-seq (stoeckius2017simultaneous)) link RNA, ATAC, and protein modalities, providing supervision for cross-view alignment. Spatial technologies (e.g. Visium  (oliveira2025high), Slide-seqV2  (stickels2021highly), MERFISH  (chen2015spatially), Stereo-seq  (chen2022spatiotemporal)) connect molecular profiles to tissue architecture, while functional datasets (e.g. Perturb-seq (replogle2022_cell)) benchmark causal reasoning. Plant resources (e.g. scPlantDB  he2024sty, Arabidopsis E-CURD-4  (shulse2019high)) extend modeling beyond animal systems, opening cross-kingdom evaluation.
@@ 4.1 Foundation Models
Foundation models learn transferable cell and gene embeddings directly from large-scale scRNA-seq without explicit labels. Representative systems such as scGPT  (cui2024scgpt), Geneformer  (theodoris2023transfer), and scFoundation  (hao2024large) pretrain on multi-tissue atlases exceeding one million cells, using masked-gene or rank-based reconstruction to capture expression context. Variants like tGPT  (shen2023generative), scBERT  (yang2022scbert), and scGraphformer  (fan2024scgraphformer) treat genes as tokens or graph nodes, while cross-species models (UCE  (rosen2023universal), GeneCompass (yang2024genecompass)) apply contrastive alignment across homologous genes. These models underpin most single-cell LLM pipelines, offering strong transfer for annotation and integration but limited ontology grounding and explainability, motivating later text-bridged and reasoning frameworks.
@@ 4.2 Text-Bridge LLMs
Text-Bridge models couple molecular embeddings with biomedical language to ground single-cell representations in semantics and ontology. Systems such as scELMo  (liu2023scelmo), CellLM  (zhao2023large), and GenePT  (chen2024genept) align gene or cell embeddings with textual descriptors, while Cell2Sentence  (pmlr-v235-levine24a) and Cell2Text  (kharouiche2025cell2text) translate expression profiles into natural-language summaries. Using dual-encoder or encoder–decoder architectures, these frameworks employ contrastive or prompt-based alignment between molecular and text encoders (e.g., BioBERT). They enhance interpretability and enable zero-shot annotation but depend heavily on curated corpora and remain non-agentic, serving as a bridge between foundation and reasoning frameworks.
@@ 4.3 Spatial and Multimodal Models
Spatial and multimodal frameworks integrate gene expression with spatial coordinates, histology, or additional omics to capture tissue architecture. Representative models include TransformerST  (lu2024sctrans), spaLLM (10778152), and OmiCLIP (cui2025towards) for spatial mapping, and scMMGPT (shi2025multimodal) and FmH2ST for RNA–ATAC–protein integration. They use multi-branch Transformers with modality-specific encoders and cross-attention fusion, trained on datasets such as Visium DLPFC and MERFISH. These models achieve strong spatial alignment and biological realism but face heterogeneous resolutions, limited open benchmarks, and high computational cost.
@@ 4.4 Epigenomic Models
Epigenomic foundation models extend LLM concepts to chromatin-accessibility and regulatory data such as scATAC-seq. EpiFoundation  (wu2025epifoundation), EpiBERT (javed2025multi), and EpiAttend (li2022epiattend) learn cis-regulatory patterns from ENCODE and other compendia, while GeneMamba and scMamba use efficient state-space layers for long-range dependency modeling. Cross-omic variants (ChromFound (jiao2025chromfound), GET (fu2025foundation)) jointly embed RNA and ATAC to infer gene-regulatory networks. These models improve biological grounding but remain constrained by sparse data and lack unified benchmarks across regulatory modalities.
@@ 4.5 Agentic Frameworks
Agentic systems integrate pretrained models with reasoning modules for autonomous single-cell analysis. Frameworks such as scAgent (mao2025scagent), CellVerse (zhang2025cellverse), and Teddy (chevalier2025teddy) combine domain-specific encoders with LLM controllers that plan tasks, query ontologies, and interface with tools or APIs. EpiAgent (chen2025epiagent) extends this paradigm to regulatory genomics. These systems enable dialogue-based annotation and multi-step reasoning but lack standardized benchmarks for reasoning fidelity and rely on underlying LLM reliability.
@@ 5.1 Annotation and Ontology Mapping
Annotation is the most common single-cell task, spanning automated cell-type labeling, ontology alignment, and cross-dataset harmonization. Foundation models such as scGPT, scFoundation, and scBERT achieve high accuracy on atlases like Tabula Sapiens and the Human Cell Atlas. Text-bridge models (CellLM  (zhao2023large), scELMo  (liu2023scelmo), GenePT  (chen2024genept)) add ontology-based interpretability, and agentic systems (scAgent) perform multi-step annotation reasoning via LLM controllers. Models are fine-tuned on curated references (e.g., Azimuth PBMC, HCA Lung) using masked-gene or contrastive objectives. Evaluation uses label accuracy or ARI; key limitations include rare-cell detection, cross-species consistency, and lack of reasoning benchmarks.
@@ 5.2 Trajectory and Perturbation Modeling
Trajectory modeling captures dynamic state transitions and causal responses to interventions. Geneformer  (theodoris2023transfer), scGPT (cui2024scgpt), and scRDiT  (dong2025scrdit) model temporal or perturbation trajectories using denoising and contrastive objectives, trained on datasets such as Replogle 2022 Perturb-seq and sci-Plex. Epigenomic models (EpiFoundation (wu2025epifoundation) ,EpiAgent (chen2025epiagent)) extend to regulatory responses through chromatin context and ontology reasoning. Metrics include expression-profile correlation and target-gene recovery. While these models predict single perturbations effectively, performance degrades for combinatorial or long-range effects, underscoring the need for unified causal benchmarks.
@@ 5.3 Multi-omic Integration
Multi-omic integration seeks unified representations across paired RNA, ATAC, and protein modalities to capture gene-regulatory and signaling relationships. Representative models include scMMGPT (shi2025multimodal), GET (General Expression Transformer) (fu2025foundation), ChromFound (jiao2025chromfound), and epigenomic-aware variants such as EpiFoundation (wu2025epifoundation) and EpiBERT (javed2025multi). These systems train on joint datasets like 10x Multiome PBMC 10k, TEA-seq, DOGMA-seq, and ASAP-seq, using cross-modal transformers or contrastive alignment between expression and accessibility features.
@@ 5.4 Spatial Mapping and Deconvolution
Spatial mapping links molecular profiles to their tissue locations, enabling inference of cell organization and microenvironmental context. Representative models include TransformerST  (lu2024sctrans), spaLLM , OmiCLIP, FmH2ST, HEIST  (madhu2025heist), and Spatial2Sentence (chen2025spatial), trained on datasets such as Visium DLPFC, Slide-seqV2, MERFISH, and Xenium. They combine spatial coordinates, histology features, and gene-expression embeddings through cross-attention or contrastive alignment to reconstruct cell or spot-level maps.
@@ 5.5 Regulatory Network and Pathway Inference
Regulatory-network and pathway inference focuses on identifying gene–gene or enhancer–promoter dependencies that drive transcriptional programs. Representative models include GeneMamba (qi2025bidirectional), scMamba (yuan2025scmamba), EpiFoundation (wu2025epifoundation), EpiBERT (javed2025multi), ChromFound (jiao2025chromfound), and GET (fu2025foundation), trained on large-scale ATAC and multiome atlases such as ENCODE, human cCRE datasets, and 10x Multiome PBMC. These models treat chromatin regions or genes as tokens and learn context-dependent regulatory relationships through self- or cross-attention.
GeneMamba (qi2025bidirectional) and scMamba employ state-space layers for efficient modeling of long-range chromatin dependencies, while EpiBERT and EpiFoundation use masked-region reconstruction to capture enhancer–promoter coupling. ChromFound and GET link RNA and accessibility signals to infer transcription-factor activity and causal directionality. Evaluation typically uses motif enrichment, AUROC for known interactions, or pathway-level overlap with KEGG or Reactome references. These models improve biological grounding and causal interpretability but remain constrained by sparse training data and lack of unified ground-truth regulatory benchmarks.
@@ 5.6 Cross-Species Translation
Cross-species translation aims to transfer cell and gene representations between organisms, enabling comparative and evolutionary analysis of single-cell atlases. Representative models include iSEEEK  (shen2022universal), UCE (Universal Cell Embeddings)  (rosen2023universal), GeneCompass  (yang2024genecompass), and scPlantLLM  (scplantlmm), trained on paired or homolog-mapped datasets such as tthe Mouse Cell Atlas, Tabula Muris, and scPlantDB. These systems learn alignment between orthologous genes or conserved cell states to support zero-shot annotation across species.
iSEEEK  (shen2022universal) integrates over 11 million human and mouse cells using shared gene-token vocabularies, while UCE  (shen2023generative) applies masked autoencoding and contrastive alignment across ortholog groups. GeneCompass  (yang2024genecompass) extends this with knowledge-informed pretraining that links species via ontology-derived embeddings, and scPlantLLM adapts the paradigm to plant single-cell data across 17 species. Evaluation uses transfer accuracy and embedding consistency across species pairs. Despite improved cross-domain generalization, these models remain limited by incomplete ortholog mappings and domain shifts in sequencing depth and tissue composition.
@@ 5.7 Generation and Simulation
Generative modeling aims to synthesize realistic single-cell profiles, simulate perturbations, or reconstruct missing modalities. Representative approaches include scGPT, scFoundation, CellFM  (zeng2025cellfm), and Geneformer, which employ decoder or diffusion-style architectures to model the probability distribution of gene-expression states. These models are pretrained on large atlases such as Tabula Sapiens and HCA, then fine-tuned to generate new cell embeddings or perturbation outcomes.
@@ 5.8 Drug-Response Prediction
Drug-response prediction models aim to infer cellular transcriptional outcomes following chemical or genetic perturbations, supporting virtual screening and mechanism discovery. Representative frameworks include Geneformer (theodoris2023transfer), scGPT  (cui2024scgpt), EpiFoundation (wu2025epifoundation), and EpiAgent (chen2025epiagent), trained or evaluated on large-scale pharmacotranscriptomic resources such as sci-Plex, Replogle 2022 Perturb-seq, and the ARC Virtual Cell Challenge. These systems learn conditional embeddings that map control to perturbed states, enabling zero-shot prediction of unseen compounds or targets.